OpenStudy (anonymous):

write an equation of the line that passes through the given points. (-5,3/4), (-2,-3/4)

OpenStudy (anonymous):

Use the two point formula of straight line: \[y-y1=((y2-y1)/(x2-x1)) \times (x-x1)\] \[y - 3/4 = (((-3/4) - 3/4)/(-2 - (-5))) \times (x-(-5))\] This is the equation which you can furthur simplify.